Project a3106 (run as a2934) (Barbara
again...) reported baseline problems with lbw polB on 26mar17
. On 28mar17 we took some data with lbw and the interim
correlator to check the receiver stability.
26mar17 a3106 data
a3106 data with polb drifts (.ps)
- lbw, linear pol, freq around 1390.
- 5 min onoffs with cal using interim correlator
- 12.5 MHz Bw, 2048 channels.
- filters used:
- 1320 hipass (lbfb 7)
- 1370 hipass with 1230-1470 bandpas (lbfb 1 3)
- Top frame: total power from 5 onoff pairs (file
- black is polA, red is polB.
- the narrow jumps are the cal.
- On the start of the 5th pair (around sample 2490) polb
started to wander around.
- Bottom frame: blowup of 5th on,off pair showing polB
The Big change from on/off pair 4 to pair
5 is that the lbw rf filters were switched:
- from 1320 hipass (lbfb 7)
- to 1370 hipass and 1230-1470 bandpass (lbfb
It looks like the filter switches are not
making good contact when switching.
28mar17 interim cor test data
On 28mar17 we took some data with lbw and
the interim correlator.
The setup was:
- lbw , linear polarization, cfr 1397 MHz
- interim correlator , 1 sec integrations, 25 MHz bandwidth.
- Telescope was being used by maintenance. It was mostly
stationary. toward the end they moved it a little.
- We setup a filter set, adjusted the power and then started
- During the scan, we switched to another set of filters
(causing a glitch) and then back to the original set. If
thing were working correctly, we should come back to the
same power levels..
The plots show the total power vs
time data during the filter switches (.ps) (.pdf)
- the data comes from the 4th scan of file
- we started with the 1370 MHz hipass and 1230-1470
band pass filters in (lbfb 1 3)
- the switch sequences flagged in purple were:
- 2190 : (1370hp,1230-1470bp) -> (1320 hp) ->
- 2730 : same sequence
- the level changed after each switch sequence for both
polA and B.
- They too 10's of seconds to approach their original
- The other jumps were similar filter switches.
- the general baseline variation is probably from the
maintenance crew moving the telescope.
We also tried switching between:
- 1230-1470bp -> 1120-1800bp -> 1230->1470bp
- this also showed a long recovery time for polb.
- It looks like the switches in the lbw filter bank are
occasionally not making good contacts (this could be the
through switch in use, or one of the 50 ohm load switches of
the unused filters).
- We need to bring the lbw rf filterbank down to the lab and
test the switches.